OpenStudy (anonymous):

Solve for B. D= 3a + 3b

OpenStudy (anonymous):

1. subtract 3a from both sides 2. divide both sides by 3 you wanna try ?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

i'm a bit confused...

OpenStudy (anonymous):

D = 3a + 3b (subtract 3a from both sides) D - 3a = 3a - 3a + 3b D - 3a = 3b (divide both sides by 3) (D - 3a)/3 = b or it could also be written like : D/3 - a = b

OpenStudy (anonymous):

do you have any questions ?

OpenStudy (anonymous):

Oh ok, thx for the help, i'll use that method from now on. Thx again!

OpenStudy (anonymous):

no problem :)